OpenStudy (anonymous):

How to find the maximum area of a triangle if given the perimeter is 10 ?

zepdrix (zepdrix):

Right triangle? <.<

OpenStudy (anonymous):

didnt say it in the question :)

zepdrix (zepdrix):

hmmmmmmmm

OpenStudy (anonymous):

i tried suppose this triangle has maximum area if it is a right triangle, but this be more complicated for me :(

zepdrix (zepdrix):

I was trying to use Heron's Formula, but I can't seem to get it in terms of a single variable. Or maybe using \(\Large\rm Area=a b\cos \theta\) where a and b are the legs around angle theta. But that introduces an angle...... doesn't seem very helpful... hmm

OpenStudy (e.mccormick):

Hmmm... I thnk you will be stuck with Heron's Formula.

OpenStudy (e.mccormick):

That, and the Isoperimetric Property of Equilateral Triangles

OpenStudy (e.mccormick):

I had to look up the name of it. LOL. I remembered that equilateral triangles had the largest area, but could not remember the name of the theorem. That simplifies Heron's Formula down a lot: \(A=\sqrt{s(s-a)(s-b)(s-c)} \text{ and } a=b=c \implies \) \(A=\sqrt{s\left(s-\dfrac{2}{3}s\right )\left(s-\dfrac{2}{3}s\right )\left(s-\dfrac{2}{3}s\right )} \implies \) \(A=\sqrt{3s\left(s-\dfrac{2}{3}s\right )} \implies \) \(A=\sqrt{3s^2-2s} \) And as always, \(s=\dfrac{a+b+c}{2}\)

OpenStudy (anonymous):

pythagoras theorem?

Miracrown (miracrown):

Think of it like a rectangle... The maximum area triangle has to be an equilateral triangle, because all the inner angles being equal optimizes the area. It's the same way a rectangle with the highest area has to be a square.

Miracrown (miracrown):

In this case we don't know the triangle is a right triangle. To maximize the area of any general shape, you want the sides and angles to be equal So this triangle is an equilateral triangle.

OpenStudy (anonymous):

okay, now be easy for mee to determine the area of this triangle if it is an equi triangle. thanks guys :)

ganeshie8 (ganeshie8):

for general case, you can set it up as an optimization problem : Maximize : \(A(a, b, c) = \sqrt{ 5(5-a)(5-b)(5-c)} \) subject to : \(a+b+c = 10\)

OpenStudy (anonymous):

can i hv a medal?

Miracrown (miracrown):

 We know all the angle measurements are the same. Since triangles add to 180 degrees we divide by 3 to find all the angles are 60 degrees Here we're looking at side measure. Each side is the same value, so I'll just say variable b equals the length of each side. Each side is b units long. What is the total perimeter?